# Bialgebras, and Lie monoid actions in Morse and Floer theory, I

### v1

@@ 3.2. $f$-bialgebras
Take $(\underline{k},\underline{l})=((k),(1))$. If $(\underline{k},\underline{l})=(\underline{k}^{1},\underline{l}^{1})\sharp(\underline{k}^{0},\underline{l}^{0})$, then one must have $\underline{l}^{0}=\underline{l}^{1}=(1)$, $\underline{k}^{0}=(k^{0})$, and condition $c(\underline{k}^{1},\underline{l}^{1})=1$ implies that $\underline{k}^{1}$ is of the form $\underline{k}^{1}=(1,\ldots,1,k^{1},1,\ldots,1)$, with $k^{1}$ at position $i$, $1\leq i\leq k^{0}$. One then has:
@@
(3.22) $v(\underline{k}^{0})$ $=k^{0}-1,$
(3.23) $v(\underline{k}^{1})$ $=k^{1}-1,$
(3.24) $c(\underline{k}^{0},\underline{l}^{0})$ $=c(\underline{k}^{1},\underline{l}^{1})=1,$
(3.25) $\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}$ $=0,$
(3.26) $\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}$ $=(k^{1}-1)(k^{0}-i).$
From Proposition 2.13, it follows that $\rho^{\underline{\underline{k}}}_{\underline{\underline{l}}}=k^{0}k^{1}+i(k^{1}+1)\text{ mod }2$. Finally, $(V^{\underline{k}^{1}}_{1})$ implies that
(3.27) $\alpha^{\underline{k}^{1}}_{1}=(-1)^{\spadesuit_{\tilde{\underline{k}}^{1}}}\cdot\left(id^{\otimes(i-1)}\otimes m^{k^{1}}\otimes id^{\otimes(k^{0}-i)}\right),$
with ${\spadesuit_{\tilde{\underline{k}}^{1}}}=0$. The $A_{\infty}$-relation (3.19) follows. The proof of the $A_{\infty}$-coalgebra statement is analogous. ∎
@@
In the first case, $\rho^{\underline{\underline{k}}}_{\underline{\underline{l}}}=1$; and in the second, $\rho^{\underline{\underline{k}}}_{\underline{\underline{l}}}=0$. Furthermore, $(D^{(1,1);1}_{(2)})$ and $(D^{(2)}_{(1,1);1})$ give respectively:
@@
Say that $(A,{\alpha}^{\underline{k}}_{\underline{l}})$ is a dg $f$-bialgebra if ${\alpha}^{\underline{k}}_{\underline{l}}=0$ unless $\deg{\alpha}^{\underline{k}}_{\underline{l}}=0$ (i.e. $\underline{k}$ is vertical and $\underline{l}$ almost-vertical, or vice versa). Then, dg $f$-bialgebras are in one-to-one correspondence with dg bialgebras.
@@
If $(A,{\alpha}^{\underline{k}}_{\underline{l}})$ is a dg $f$-bialgebra, then it follows from $(R^{2}_{1})$, $(R^{3}_{1})$, $(R^{1}_{2})$, $(R^{1}_{3})$ and $(R^{2}_{2})$ that $(A,\partial,\alpha^{2}_{1},\alpha^{1}_{2})$ is a dg bialgebra. Conversely, given a dg bialgebra, relations $(V)$ and $(D)$ prescribe the degree zero ${\alpha}^{\underline{k}}_{\underline{l}}$ maps. Setting the degree nonzero operations equal to zero gives a dg $f$-bialgebra.
@@ 3.3. Morphisms of $f$-bialgebras
Just like the coherence relations of $A_{\infty}$-morphisms come from the boundary structure of the multiplihedron, our notion of morphisms between $f$-bialgebras comes from the boundary structure of $J^{\underline{k}}_{\underline{l}}$. Let us split:
(3.35) $GBF_{1}(\underline{k},\underline{l})=GBF_{1}^{f=0}(\underline{k},\underline{l})\cup GBF_{1}^{f=1}(\underline{k},\underline{l}).$
@@
($M^{\underline{k}}_{\underline{l}}$) $\partial(f^{\underline{k}}_{\underline{l}})=\sum_{GBF_{1}^{f=0}(\underline{k},\underline{l})}(-1)^{\rho^{\underline{\underline{k}}}_{\underline{\underline{l}};0}}\cdot f^{\underline{k}^{0}}_{\underline{l}^{0}}\circ{\alpha}^{\underline{k}^{1}}_{\underline{l}^{1}}+\sum_{GBF_{1}^{f=1}(\underline{k},\underline{l})}(-1)^{\rho^{\underline{\underline{k}}}_{\underline{\underline{l}};1}}\cdot{\beta}^{\underline{k}^{0}}_{\underline{l}^{0}}\circ f^{\underline{k}^{1}}_{\underline{l}^{1}},$
where we recall that $GBF_{1}^{f=0}(\underline{k},\underline{l})$ and $GBF_{1}^{f=1}(\underline{k},\underline{l})$ are described explicitly in (2.60) and (2.61). Furthermore $f$ satisfies the simplification relations:
@@
    ($W^{\underline{k}}_{b}$) $f^{\underline{k}}_{\underline{l}}\simeq f^{k_{1}}_{\underline{l}}\otimes\cdots\otimes f^{k_{a}}_{\underline{l}},$
    and
    ($X^{1}_{b}$) $f^{1}_{\underline{l}}\simeq f^{1}_{1}\otimes\cdots\otimes f^{1}_{1}.$
@@
    and
    ($X^{a}_{1}$) $f^{\underline{k}}_{1}\simeq f^{1}_{1}\otimes\cdots\otimes f^{1}_{1}.$
@@
Consider $f\colon A\to B$ as in Definition 3.6 above. Then the family $\varphi^{k}=f^{k}_{1}$ defines an $A_{\infty}$-morphism from $A$ to $B$, endowed with the $A_{\infty}$-algebra structures $m_{A}^{k}$, $m_{B}^{k}$ of Proposition 3.4, namely they satisfy:
(3.37) $\partial(\varphi^{k})$ $=\sum_{\begin{subarray}{c}k+1=k^{0}+k^{1},\\ 1\leq i\leq k^{0}\end{subarray}}(-1)^{k^{1}k^{0}+(k^{1}+1)i}\cdot\varphi^{k^{0}}\circ\left(id^{\otimes(i-1)}\otimes m_{A}^{k^{1}}\otimes id^{\otimes(k^{0}-i)}\right)$
$+\sum_{\begin{subarray}{c}\left|\underline{k}^{1}\right|=k,\\ n(\underline{k}^{1})=k^{0}\end{subarray}}(-1)^{\sum_{i=1}^{k^{0}}(k^{1}_{i}+1)(k^{0}+i)+1}\cdot m_{B}^{k^{0}}\circ\left(\varphi^{k^{1}_{1}}\otimes\cdots\otimes\varphi^{k^{1}_{k^{0}}}\right).$
@@
(3.38) $\partial(\psi^{l})$ $=\sum_{\begin{subarray}{c}l+1=l^{0}+l^{1},\\ 1\leq j\leq l^{1}\end{subarray}}(-1)^{(l^{0}+1)(j+1)+1}\cdot\left(id^{\otimes(j-1)}\otimes\delta_{B}^{l^{0}}\otimes id^{\otimes(l^{1}-j)}\right)\circ\psi^{l^{1}}$
$+\sum_{\begin{subarray}{c}\left|\underline{l}^{0}\right|=l,\\ n(\underline{l}^{0})=l^{1}\end{subarray}}(-1)^{\sum_{j=1}^{l^{1}}j(l^{0}_{j}+1)+1}\cdot\left(\psi^{l^{0}_{1}}\otimes\cdots\otimes\psi^{l^{0}_{l^{1}}}\right)\circ\delta_{A}^{l^{1}}.$
@@
Let us first consider $GBF^{f=0}_{1}(k,1)$: $\underline{l}^{0}=\underline{l}^{1}=(1)$, $\underline{k}^{0}=(k^{0})$, and $c(\underline{k}^{1},\underline{l}^{1})=1$ implies that $\underline{k}^{1}$ is of the form $\underline{k}^{1}=(1,\ldots,1,k^{1},1,\ldots,1)$, with $k^{1}$ at position $i$, $1\leq i\leq k^{0}$. One then has:
@@
(3.40) $v(\underline{k}^{0})$ $=k^{0}-1,$
(3.41) $v(\underline{k}^{1})$ $=k^{1}-1,$
(3.42) $c(\underline{k}^{0},\underline{l}^{0})$ $=c(\underline{k}^{1},\underline{l}^{1})=1,$
(3.43) $\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}$ $=0,$
(3.44) $\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}$ $=(k^{1}-1)(k^{0}-i).$
From Proposition 2.13, it follows that $\rho^{\underline{\underline{k}}}_{\underline{\underline{l}};0}=k^{0}k^{1}+i(k^{1}+1)\text{ mod }2$. Finally, $(V^{\underline{k}^{1}}_{1})$ gives
(3.45) $\alpha^{\underline{k}^{1}}_{1}=(-1)^{\spadesuit_{\tilde{\underline{k}}^{1}}}\cdot\left(id^{\otimes(i-1)}\otimes m_{A}^{k^{1}}\otimes id^{\otimes(k^{0}-i)}\right),$
with ${\spadesuit_{\tilde{\underline{k}}^{1}}}=0$.
Consider now $GBF^{f=1}_{1}(k,1)$: as for $f=0$, $\underline{l}^{0}=\underline{l}^{1}=(1)$, $\underline{k}^{0}=(k^{0})$, but now there are no constraints on $\underline{k}^{1}$, other than $\left|\underline{k}^{1}\right|=k$ and $n(\underline{k}^{1})=k^{0}$. One has:
@@
(3.47) $v(\underline{k}^{0})$ $=k^{0}-1,$
(3.48) $v(\underline{k}^{1})$ $=\sum_{i=1}^{k^{0}}(k^{1}_{i}-1),$
(3.49) $c(\underline{k}^{0},\underline{l}^{0})$ $=c(\underline{k}^{1},\underline{l}^{1})=1,$
(3.50) $\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}$ $=0,$
(3.51) $\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}$ $=\sum_{i=1}^{k^{0}}(k^{1}_{i}-1)(k^{0}-i),$
@@
with the sign given by the orientation of the boundary strata of $J(2)^{\underline{k}}_{\underline{l}}$. With this operation, $f$-bialgebras form a category. However, mapping a Lie group to its Morse complex does not define a strict functor: $(g\circ f)_{*}=g_{*}\circ f_{*}$ only holds up to a homotopy, measured by $J(2)$. It should instead be an $\infty$-functor: $f\text{-}\mathcal{B}ialg$ should be enhanced to an $(\infty,1)$-category, using the higher multiplihedron $J(n)$. We postpone the details of the higher categorical structure of $f\text{-}\mathcal{B}ialg$ to later.
@@ 3.4. Bimodules of $f$-bialgebras
Define $BF_{1}(\underline{k}^{l}|1|\underline{k}^{r},\underline{l})$ from $BF_{1}(\underline{k},\underline{l})$ by splitting each floor $f=0,1$ multi-index $\underline{k}_{f}$ to $\underline{k}_{f}^{l}|1|\underline{k}_{f}^{r}$ at the appropriate position, see Figure 15, i.e. we first split $\underline{k}_{1}$ at position $i_{1}=\left|\underline{k}^{l}\right|+1$ to get $\underline{k}_{1}^{l}|1|\underline{k}_{1}^{r}$, then split $\underline{k}_{0}$ at position $i_{0}=n(\underline{k}_{1}^{l})$ to get $\underline{k}_{0}^{l}|1|\underline{k}_{0}^{r}$.
@@
If now $\underline{k}^{l},\underline{k}^{r}$ are multi-indices with only entries $\geq 1$, we can define $BF_{1}(\underline{k}^{l}|0|\underline{k}^{r},\underline{l})$ from $BF_{1}(\underline{k},\underline{l})$ similarly, with now
@@
Let $(A,\partial,\left\{{\alpha}^{\underline{k}}_{\underline{l}}\right\})$ and $(B,\partial,\left\{{\beta}^{\underline{k}}_{\underline{l}}\right\})$ be two $f$-bialgebras. An $(A,B)$ pre $u$-bimodule (of $f$-bialgebras) is a chain complex $(M,\partial)$, with a collection of operations, with $\epsilon=0,1$:
@@
satisfying the family of relations:
($R^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}$) $\partial(\mu^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}})=\sum_{BF_{1}(\underline{k}^{l}|\epsilon|\underline{k}^{r},\underline{l})}(-1)^{\rho^{\underline{\underline{k}}}_{\underline{\underline{l}}}}\cdot\mu^{\underline{k}^{l}_{0}|\epsilon|\underline{k}^{r}_{0}}_{\underline{l}_{0}}\circ\mu^{\underline{k}^{l}_{1}|\epsilon|\underline{k}^{r}_{1}}_{\underline{l}_{1}}.$
We will say that $M$ is an $(A,B)$ $u$-bimodule if furthermore:
  • $\mu^{0|1|0}_{1}=id_{M}$,
@@
  • (vertical forest splittings) If $\underline{l}=(1,\ldots,1)$, then
    ($V^{\underline{k}^{l}|1|\underline{k}^{r}}_{b}$) $\mu^{\underline{k}^{l}|1|\underline{k}^{r}}_{\underline{l}}$ $\simeq(-1)^{\spadesuit_{\tilde{k}}}\cdot{\alpha}^{k_{1}^{l}}_{\underline{l}}\otimes\cdots\otimes\mu^{k^{l}_{a^{l}+1}|1|k^{r}_{0}}_{\underline{l}}\otimes\cdots\otimes{\beta}^{k_{a^{r}}^{r}}_{\underline{l}},$
    ($V^{\underline{k}^{l}|0|\underline{k}^{r}}_{b}$) $\mu^{\underline{k}^{l}|0|\underline{k}^{r}}_{\underline{l}}$ $\simeq(-1)^{\spadesuit_{\tilde{k}}}\cdot{\alpha}^{k_{1}^{l}}_{\underline{l}}\otimes\cdots\otimes{\alpha}^{k_{a^{l}}^{l}}_{\underline{l}}\otimes{\beta}^{k_{1}^{r}}_{\underline{l}}\otimes\cdots\otimes{\beta}^{k_{a^{r}}^{r}}_{\underline{l}}.$
    Likewise, if $\underline{k}=(1,\ldots,1)$, i.e. either $\epsilon=1$ and $\underline{k}^{l}=(1,\ldots,1,0)$ and $\underline{k}^{r}=(0,1,\ldots,1)$, or $\epsilon=0$ and $\underline{k}^{l}=(1,\ldots,1)$ and $\underline{k}^{r}=(1,\ldots,1)$, then
    ($V^{a^{l}|\epsilon|a^{r}}_{\underline{l}}$) $\mu^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}\simeq(-1)^{\spadesuit_{\tilde{l}}}\cdot\mu^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{l_{1}}\otimes\cdots\otimes\mu^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{l_{b}}.$

### v2

@@ 3.2. $f$-bialgebras
Take $\underline{k}=(k)$ and $\underline{l}=(1)$. If $\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}$ and $\underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}$, then one must have $\underline{l}^{0}=\underline{l}^{1}=(1)$, $\underline{k}^{0}=(k^{0})$, and condition $c(\underline{k}^{1},\underline{l}^{1})=1$ implies that $\underline{k}^{1}$ is of the form $\underline{k}^{1}=(1,\ldots,1,k^{1},1,\ldots,1)$, with $k^{1}$ at position $i$, $1\leq i\leq k^{0}$. One then has:
@@
(3.19) $v(\underline{k}^{0})$ $=k^{0}-1,\ v(\underline{k}^{1})=k^{1}-1,$
(3.20) $\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}$ $=0,\ \heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}=(k^{1}-1)(k^{0}-i).$
From Proposition 2.12, it follows that $\rho=k^{0}k^{1}+i(k^{1}+1)\text{ mod }2$. Finally, $(V^{\underline{k}^{1}}_{1})$ implies that
(3.21) $\alpha^{\underline{k}^{1}}_{1}=\left(id^{\otimes(i-1)}\otimes m^{k^{1}}\otimes id^{\otimes(k^{0}-i)}\right).$
The $A_{\infty}$-relation (3.16) follows. The proof of the $A_{\infty}$-coalgebra statement is analogous. ∎
@@
In the first case, $\rho=1$; and in the second, $\rho=0$. Furthermore, $(D^{(1,1);1}_{(2)})$ and $(D^{(2)}_{(1,1);1})$ give respectively:
@@
Say that $(A,{\alpha}^{\underline{k}}_{\underline{l}})$ is a dg $f$-bialgebra if ${\alpha}^{\underline{k}}_{\underline{l}}=0$ unless $\deg{\alpha}^{\underline{k}}_{\underline{l}}\leq 0$. Then, dg $f$-bialgebras are in one-to-one correspondence with dg bialgebras.
@@
If $(A,{\alpha}^{\underline{k}}_{\underline{l}})$ is a dg $f$-bialgebra, then it follows from $(R^{2}_{1})$, $(R^{3}_{1})$, $(R^{1}_{2})$, $(R^{1}_{3})$ and $(R^{2}_{2})$ that $(A,\partial,\alpha^{2}_{1},\alpha^{1}_{2})$ is a dg bialgebra. Conversely, given a dg bialgebra, relations $(V)$ and $(D)$ prescribe the degree $\leq 0$ maps. Setting the positive degree operations equal to zero gives a dg $f$-bialgebra.
In many cases, some maps $\alpha^{\underline{k}}_{\underline{l}}$ must vanish for degree reasons:
Proposition 3.8.
If $A$ is $\mathbb{Z}$-graded and is concentrated in degrees 0, …, $\mathrm{maxdeg}A$, then $\alpha^{\underline{k}}_{\underline{l}}=0$ if $\left|\underline{k}\right|>(a\cdot\mathrm{maxdeg}A-1)\left|\underline{l}\right|+a+b+1.$
Proof.
${\alpha}^{\underline{k}}_{\underline{l}}\colon(A^{b})^{\left|\underline{k}\right|}\to(A^{\left|\underline{l}\right|})^{a}$ has degree $\deg{\alpha}^{\underline{k}}_{\underline{l}}=v(\underline{k})+v(\underline{l})-1=\left|\underline{k}\right|-a+\left|\underline{l}\right|-b-1$, therefore it is zero if
(3.29) $\deg{\alpha}^{\underline{k}}_{\underline{l}}+\mathrm{mindeg}(A^{b})^{\left|\underline{k}\right|}>\mathrm{maxdeg}(A^{\left|\underline{l}\right|})^{a}.$
Furthermore, $\mathrm{mindeg}(A^{b})^{\left|\underline{k}\right|}\geq 0$ and $\mathrm{maxdeg}(A^{\left|\underline{l}\right|})^{a}=a\left|\underline{l}\right|\mathrm{maxdeg}A$. The claim follows. ∎
@@ 3.3. Morphisms of $f$-bialgebras
Just like the coherence relations of $A_{\infty}$-morphisms come from the boundary structure of the multiplihedron, our notion of morphisms between $f$-bialgebras comes from the boundary structure of $J^{\underline{k}}_{\underline{l}}$.
@@
($M^{\underline{k}}_{\underline{l}}$) $0=\sum_{\begin{subarray}{c}\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}\\ \underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}\end{subarray}}(-1)^{\rho_{0}}\cdot f^{\underline{k}^{0}}_{\underline{l}^{0}}\circ{\alpha}^{\underline{k}^{1}}_{\underline{l}^{1}}+(-1)^{\rho_{1}}\cdot{\beta}^{\underline{k}^{0}}_{\underline{l}^{0}}\circ f^{\underline{k}^{1}}_{\underline{l}^{1}},$
Furthermore, $f$ satisfies the simplification relations:
@@
    ($W^{\underline{k}}_{b}$) $f^{\underline{k}}_{\underline{l}}\simeq f^{k_{1}}_{\underline{l}}\otimes\cdots\otimes f^{k_{a}}_{\underline{l}}.$
@@
Remark 3.10.
Since ${\alpha}^{\underline{k}^{1}}_{\underline{l}^{1}}=0$ if $c(\underline{k}^{1},\underline{l}^{1})\geq 2$ and ${\beta}^{\underline{k}^{0}}_{\underline{l}^{0}}=0$ if $c(\underline{k}^{0},\underline{l}^{0})\geq 2$, one can write ($M^{\underline{k}}_{\underline{l}}$) as:
(3.31) $\partial(f^{\underline{k}}_{\underline{l}})=\sum_{c(\underline{k}^{1},\underline{l}^{1})=1}(-1)^{\rho_{0}}\cdot f^{\underline{k}^{0}}_{\underline{l}^{0}}\circ{\alpha}^{\underline{k}^{1}}_{\underline{l}^{1}}+\sum_{c(\underline{k}^{0},\underline{l}^{0})=1}(-1)^{\rho_{1}}\cdot{\beta}^{\underline{k}^{0}}_{\underline{l}^{0}}\circ f^{\underline{k}^{1}}_{\underline{l}^{1}}.$
@@
Consider $f\colon A\to B$ as in Definition 3.9 above. Then the family $\varphi^{k}=f^{k}_{1}$ defines an $A_{\infty}$-morphism from $A$ to $B$ (in the sign convention below), endowed with the $A_{\infty}$-algebra structures $m_{A}^{k}$, $m_{B}^{k}$ of Proposition 3.5, namely they satisfy:
(3.32) 00 0 $=\sum_{\begin{subarray}{c}k+1=k^{0}+k^{1},\\ 1\leq i\leq k^{0}\end{subarray}}(-1)^{k^{1}k^{0}+(k^{1}+1)i+1}\cdot\varphi^{k^{0}}\circ\left(id^{\otimes(i-1)}\otimes m_{A}^{k^{1}}\otimes id^{\otimes(k^{0}-i)}\right)$
$+\sum_{\begin{subarray}{c}\left|\underline{k}^{1}\right|=k,\\ n(\underline{k}^{1})=k^{0}\end{subarray}}(-1)^{1+\sum_{i=1}^{k^{0}}(k^{1}_{i}+1)(k^{0}+i)}\cdot m_{B}^{k^{0}}\circ\left(\varphi^{k^{1}_{1}}\otimes\cdots\otimes\varphi^{k^{1}_{k^{0}}}\right).$
@@
(3.33) 00 0 $=\sum_{\begin{subarray}{c}l+1=l^{0}+l^{1},\\ 1\leq j\leq l^{1}\end{subarray}}(-1)^{l^{0}j+l^{0}+j}\cdot\left(id^{\otimes(j-1)}\otimes\delta_{B}^{l^{0}}\otimes id^{\otimes(l^{1}-j)}\right)\circ\psi^{l^{1}}$
$+\sum_{\begin{subarray}{c}\left|\underline{l}^{0}\right|=l,\\ n(\underline{l}^{0})=l^{1}\end{subarray}}(-1)^{\sum_{j=1}^{l^{1}}j(l^{0}_{j}+1)}\cdot\left(\psi^{l^{0}_{1}}\otimes\cdots\otimes\psi^{l^{0}_{l^{1}}}\right)\circ\delta_{A}^{l^{1}}.$
@@
From the axioms, ${\alpha}^{\underline{k}^{1}}_{\underline{l}^{1}}=0$ unless $c(\underline{k}^{1},\underline{l}^{1})\leq 1$. Here, $\underline{l}=(1)$, which implies that $\underline{k}^{1}$ is of the form $\underline{k}^{1}=(1,\ldots,1,k^{1},1,\ldots,1)$, with $k^{1}$ at position $i$, $1\leq i\leq k^{0}$. One then has:
@@
(3.35) $v(\underline{k}^{0})$ $=k^{0}-1,\ v(\underline{k}^{1})=k^{1}-1,$
(3.36) $\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}$ $=0,\ \heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}=(k^{1}-1)(k^{0}-i).$
From Proposition 2.12, it follows that $\rho_{0}=k^{0}k^{1}+i(k^{1}+1)+1\text{ mod }2$. Finally, $(V^{\underline{k}^{1}}_{1})$ gives
(3.37) $\alpha^{\underline{k}^{1}}_{1}=\left(id^{\otimes(i-1)}\otimes m_{A}^{k^{1}}\otimes id^{\otimes(k^{0}-i)}\right).$
Consider now the second sum. As for the first one, $\underline{l}^{0}=\underline{l}^{1}=(1)$, $\underline{k}^{0}=(k^{0})$, but now there are no constraints on $\underline{k}^{1}$, other than $\left|\underline{k}^{1}\right|=k$ and $n(\underline{k}^{1})=k^{0}$. One has:
@@
(3.39) $v(\underline{k}^{0})$ $=k^{0}-1,\ v(\underline{k}^{1})=\sum_{i=1}^{k^{0}}(k^{1}_{i}-1),$
(3.40) $\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}$ $=0,\ \heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}=\sum_{i=1}^{k^{0}}(k^{1}_{i}-1)(k^{0}-i),$
@@
with the sign given by the orientation of the boundary strata of $J(2)^{\underline{k}}_{\underline{l}}$. With this operation, $f$-bialgebras form a category. However, mapping a Lie group to its Morse complex does not define a strict functor: $(g\circ f)_{*}=g_{*}\circ f_{*}$ only holds up to a homotopy, measured by $J(2)$. In [Cazc], the first author will show that $f\text{-}\mathcal{B}ialg$ can be enhanced to an $(\infty,1)$-category, using the higher multiplihedron $J(n)$. With this structure, the above assignment indeed defines an $\infty$-functor.
@@ 3.4. Bimodules of $f$-bialgebras
Write $(\underline{k}^{l}|1|\underline{k}^{r})=(\underline{k}^{l}_{1}|1|\underline{k}^{r}_{1})\sharp(\underline{k}^{l}_{0}|1|\underline{k}^{r}_{0})$ if the associated total multi-indices satisfy $\underline{k}=\underline{k}_{1}\sharp\underline{k}_{0}$, and the splitting positions $i_{0},i_{1}$ are as follows (see Figure 12): we first split $\underline{k}_{1}$ at position $i_{1}=\left|\underline{k}^{l}\right|+1$ to get $\underline{k}_{1}^{l}|1|\underline{k}_{1}^{r}$, then split $\underline{k}_{0}$ at position $i_{0}=n(\underline{k}_{1}^{l})$ to get $\underline{k}_{0}^{l}|1|\underline{k}_{0}^{r}$.
We will write $(\underline{k}^{l}|0|\underline{k}^{r})$ if $\underline{k}^{l},\underline{k}^{r}$ are multi-indices with only entries $\geq 1$, and associate the total multi-index
We will write $(\underline{k}^{l}|0|\underline{k}^{r})=(\underline{k}^{l}_{1}|0|\underline{k}^{r}_{1})\sharp(\underline{k}^{l}_{0}|0|\underline{k}^{r}_{0})$ if $\underline{k}^{l}=\underline{k}^{l}_{1}\sharp\underline{k}^{l}_{0}$ and $\underline{k}^{r}=\underline{k}^{r}_{1}\sharp\underline{k}^{r}_{0}$.
@@
Let $(A,\left\{{\alpha}^{\underline{k}}_{\underline{l}}\right\})$ and $(B,\left\{{\beta}^{\underline{k}}_{\underline{l}}\right\})$ be two $f$-bialgebras. An $(A,B)$ $u$-bimodule (of $f$-bialgebras) is a chain complex $(M,\partial_{M})$, with a collection of operations, with $\epsilon=0,1$:
@@
of degree $\deg\mu^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}=v(\underline{k})+v(\underline{l})-1$, with $\underline{k}$ the total multi-index corresponding to $(\underline{k}^{l}|\epsilon|\underline{k}^{r})$, such that:
  • they satisfy the family of relations:
    ($R^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}$) $0=\sum(-1)^{\rho}\cdot\mu^{\underline{k}^{l}_{0}|\epsilon|\underline{k}^{r}_{0}}_{\underline{l}_{0}}\circ\mu^{\underline{k}^{l}_{1}|\epsilon|\underline{k}^{r}_{1}}_{\underline{l}_{1}},$
    where the sum runs over $(\underline{k}^{l}|\epsilon|\underline{k}^{r},\underline{l})=(\underline{k}^{l}_{1}|\epsilon|\underline{k}^{r}_{1})\sharp(\underline{k}^{l}_{0}|\epsilon|\underline{k}^{r}_{0})$ and $\underline{l}=\underline{l}_{0}\sharp\underline{l}_{1}$.
  • If $\underline{k}=\underline{1}_{a}$ and $\underline{l}=\underline{1}_{b}$, then $\mu^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}$ is the tensor differential induced by $\alpha^{1}_{1}$, $\partial_{M}=\mu^{0|1|0}_{1}$ and $\beta^{1}_{1}$. We will refer to this condition as $\left(T^{a^{l}|\epsilon|a^{r}}_{b}\right)$.
  • If $\underline{k}^{l}|\epsilon|\underline{k}^{r}$ is such that $\underline{k}=\underline{1}_{a}$ and $\underline{l}=(1,\ldots 1,l_{j},1,\ldots,1)$, with $l_{j}\geq 2$, then
    ($V^{a^{l}|\epsilon|a^{r}}_{\underline{l}}$) ${\mu}^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}\simeq id\otimes{\mu}^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{l_{j}}\otimes id,$
    where $\simeq$ is in the sense of (3.4), and $id$ refer to the identities of the appropriate tensor products of $A,M$ and $B$.
  • If $\underline{l}=\underline{1}_{b}$, then
    ($V^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{b}$) ${\mu}^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}\simeq\begin{cases}id\otimes{\alpha}^{k_{i}^{l}}_{\underline{l}}\otimes id,&\text{ if }\underline{k}^{l}=(1,\ldots,1,k_{i}^{l},1,\ldots,1);\ \underline{k}^{r}=\underline{1},\\ id\otimes{\mu}^{k_{a^{l}+1}^{l}|1|k_{0}^{r}}_{\underline{l}}\otimes id,&\text{if }(\underline{k}^{l}|\epsilon|\underline{k}^{r})=(1,\ldots,1,k_{a^{l}+1}^{l}|1|k_{0}^{r},1,\ldots,1),\\ id\otimes{\beta}^{k_{i}^{r}}_{\underline{l}}\otimes id,&\text{ if }\underline{k}^{l}=\underline{1};\ \underline{k}^{r}=(1,\ldots,1,k_{i}^{r},1,\ldots,1),\end{cases}$
    where in each case, $k_{i}^{l}$, $k_{a^{l}+1}^{l}+1+k_{0}^{r}$ and $k_{i}^{r}$ respectively are $\geq 2$; $\simeq$ is in the sense of (3.3); and $id$ refer to the identities of the appropriate tensor products of $A,M$ and $B$.
  • If $(\underline{k}^{l}|\epsilon|\underline{k}^{r})$ and $\underline{l}$ are such that $c(\underline{k},\underline{l})\geq 2$, then ${\mu}^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}=0$. We refer to this condition either as $(V^{a^{l}|\epsilon|a^{r}}_{\underline{l}})$ or $(V^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{b})$.